OpenStudy (anonymous):

How many rational roots does x4 + 3x3 + 3x2 + 3x + 2 = 0 have? 0 1 2 4

OpenStudy (e.cociuba):

x^4+3x^2+2=0 let x^2 = t so t^2 +3t^2 +2 =0 (t +2)(t+1)=0 so x^2 =-2 and x^2 =-1 so the equation has 4 complex roots and no rational roots. ANSWER: 4

OpenStudy (anonymous):

wow Thank YOU lol... Really appreciate it :)

OpenStudy (e.cociuba):

np:)

OpenStudy (anonymous):

Fan :)

OpenStudy (e.cociuba):

tnxs:)

OpenStudy (anonymous):

@e.cociuba did u get tht off the internet???

OpenStudy (e.cociuba):

I referenced it, to see if i got it rite:) Did u get the same answer as me?

OpenStudy (anonymous):

no, i got 6 i think u did tht the wrong way...

OpenStudy (anonymous):

i mean 2

OpenStudy (anonymous):

sry

OpenStudy (e.cociuba):

um 6 is not even a choice!! u obviously got it worng :)

OpenStudy (e.cociuba):

yea prove ur work!

OpenStudy (anonymous):

yeah i did a mistake... geez

OpenStudy (anonymous):

x^4+3x^2+2=0 let x^2 = t so t^2 +3t^2 +2 =0 (t +2)(t+1)=0 so x^2 =-2 and x^2 =-1 so the equation has 2 complex roots and no rational roots. answer= 2

OpenStudy (e.cociuba):

Tht makes no sence! You just changed my answer from 2 to 4 'honey'! Don't joke around with this prob. If u want to help, help the rite way! :D

OpenStudy (e.cociuba):

*sense

OpenStudy (anonymous):

ok im soo sry i luv u soo much.. im so sry, do u for give me??

OpenStudy (e.cociuba):

um ur forgiven.